OpenStudy (anonymous):

What is the value of 5p1?

jimthompson5910 (jim_thompson5910):

Hint: \[\Large _nP_r = \frac{n!}{(n-r)!}\]

OpenStudy (anonymous):

wait what's n and r??

jimthompson5910 (jim_thompson5910):

that's the general formula

jimthompson5910 (jim_thompson5910):

in this case, n = 5 and r = 1

OpenStudy (anonymous):

oh okay thank you!!

jimthompson5910 (jim_thompson5910):

so \[\Large _nP_r = \frac{n!}{(n-r)!}\] becomes \[\Large _5P_1 = \frac{5!}{(5-1)!}\]

OpenStudy (anonymous):

Thank you but what does the exclanation ! mean how do you figure that out?

jimthompson5910 (jim_thompson5910):

oh sry, i should have clarified

jimthompson5910 (jim_thompson5910):

n! = n*(n-1)*...3*2*1 So 5! = 5*4*3*2*1

jimthompson5910 (jim_thompson5910):

Basically you start with the given number and you multiply it by every integer smaller than it til you get to 1

OpenStudy (anonymous):

Ok thanks!

jimthompson5910 (jim_thompson5910):

you're welcome